OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x x = 10 x = 4 x = −14 x = −17

OpenStudy (anonymous):

\[\frac{ (x+20) }{ 2 }=3x\]\[(2)\frac{ (x+20) }{ 2 }=3x(2)\]\[(x+20)=6x\]\[20=6x-x\]\[20=5x\]\[\frac{ 20 }{ 5 }=x\]\[4=x\]

OpenStudy (anonymous):

@babygirlyaz123